OpenStudy (anonymous):

Having a problem with function operations. Question says Let f(x)=3x+5 and g(x)=x2. Perform each operation. The actual problem says g (x) -f(x)?

OpenStudy (anonymous):

You just have to subtract each of the functions:\[g(x)-f(x)=x^2-(3x+5)=x^2-3x-5\]and that's it!